# Differentially private and decentralized randomized power method

### v2

@@ 2.4 Existing convergence bounds
To our knowledge, the strongest bound for the randomized power method as defined in Algorithm 1 with $\Delta_{l}=\Delta_{l}^{prior}\triangleq\sqrt{p}\cdot\|\bm{X}_{l}\|_{\max}$ is given in Balcan et al. (2016). In Theorem 2.4, we present a reduction of their Theorem 3.1. The theorem, and several others we present relies on the verification of the conditions in Assumption 2.3.
@@
Let $\bm{A}\in\mathbb{R}^{d\times d}$ be a symmetric data matrix. Fix target rank $k$, intermediate rank $q\geq k$, and iteration rank $p$, with $k\leq q\leq p$. Let $\bm{U}_{q}\in\mathbb{R}^{d\times q}$ be the top-$q$ eigenvectors of $\bm{A}$ and let $\sigma_{1}\geq\cdots\geq\sigma_{d}\geq 0$ denote its eigenvalues. Let us fix:
$\eta=O\left(\frac{\sigma_{q}}{\sigma_{k}}\cdot\min\left\{\frac{1}{\log\left(\frac{\sigma_{k}}{\sigma_{q}}\right)},\frac{1}{\log\left(\tau d\right)}\right\}\right).$ (8)
@@
$\|\bm{G}_{\ell}\|_{2}$ $=O\left(\eta(\sigma_{k}-\sigma_{q+1})\right)\,,\quad\text{and}$ (9)
$\quad\|\bm{U}_{q}^{\top}\bm{G}_{\ell}\|_{2}$ $=O\left(\eta\left(\sigma_{k}-\sigma_{q+1}\right)\frac{\sqrt{p}-\sqrt{q-1}}{\tau\sqrt{d}}\right)\,.$ (10)
Theorem 2.4 (Private Power Method (PPM) (Balcan et al., 2016)).
Let $\bm{A}\in\mathbb{R}^{d\times d}$ be a symmetric data matrix. Fix target rank $k$, intermediate rank $q\geq k$, and iteration rank $p$, with $2q\leq p\leq d$. Suppose the number of iterations $L$ is set as $L=\Theta(\frac{\sigma_{k}}{\sigma_{k}-\sigma_{q+1}}\log(d))$. Let $\epsilon,\delta\in(0,1)$ be the differential privacy parameters. Let $\bm{U}_{q}\in\mathbb{R}^{d\times q}$ be the top-$q$ eigenvectors of $\bm{A}$ and let $\sigma_{1}\geq\cdots\geq\sigma_{d}\geq 0$ denote its eigenvalues. Then if Assumption 2.3 holds, Algorithm 1 with $\Delta_{l}=\Delta_{l}^{prior}$ gives us the following runtime-dependent convergence bound:
$\eta=O\left(\frac{\epsilon^{-1}\max_{l}{(\|\bm{X}_{l}\|_{\max})}\sqrt{pLd\log(1/\delta)\log(L)}}{\sigma_{k}-\sigma_{q+1}}\right)$
and also the following runtime-independent convergence bound
$\eta=O\left(\frac{\epsilon^{-1}\|\bm{U}\|_{\max}\sqrt{pLd\log(1/\delta)\log(d)\log(L)}}{\sigma_{k}-\sigma_{q+1}}\right).$
3 Analysis of Runtime-dependent Bounds
3.1 Proposed runtime-dependent bounds
3.1.1 Adjacency notion:
In prior works (Hardt & Price, 2014; Balcan et al., 2016), two matrices $\bm{A}$ and $\bm{A}^{\prime}$ are considered adjacent (denoted $\bm{A}\sim\bm{A}^{\prime}$) if they differ by a change in a single element, with a Frobenius norm of at most 1:
@@ 3.1.1 Adjacency notion:
where $\bm{e}_{i},\bm{e}_{j}\in\mathbb{R}^{n}$ are canonical basis vectors of dimension $d$, $c\leq 1$ represents the magnitude of the change, and $0\leq i,j<n$. This adjacency notion models a sensitive change as a modification to a single element in $\bm{A}$ to protect individual element-wise updates under Differential Privacy. However, since $\bm{A}$ is symmetric positive semi-definite and changes must preserve this property, this necessitates updates on the diagonal elements. We propose a new, more general, notion of adjacency to allow for other types of updates:
@@
where $\bm{C}\in\mathbb{R}^{n\times n}$ is a symmetric matrix representing the update, subject to the condition:
$\sqrt{\sum_{i=1}^{n}\|\bm{C}_{i,:}\|_{1}^{2}}\leq 1.$ (13)
The proposed adjacency notion is strictly more general than Equation (11). For example, setting $\bm{C}=c\cdot\bm{e}_{i}\bm{e}_{j}^{\top}$ recovers the original definition. Additionally, $\bm{C}$ can have non-zero entries on the diagonal, anti-diagonal, or any symmetric pattern, allowing for a variety of updates while maintaining symmetry.
In the context of recommender systems where $\bm{A}=\bm{R}^{\top}\bm{R}$ and $\bm{R}$ is a user-item interaction matrix, this adjacency notion allows for element-wise changes in $\bm{R}$ (i.e. protect user-item interactions). By adopting this more general adjacency definition, we make our privacy guarantees applicable to a wider range of real-world scenarios.
3.1.2 Improved sensitivity bound:
The previous sensitivity bound $\Delta_{l}^{\text{prior}}$ (Hardt & Price, 2014; Balcan et al., 2016) for $\bm{A}\bm{X}^{l-1}$ was estimated by extrapolating from the case where $\bm{X}^{l-1}\in\mathbb{R}^{n\times 1}$ to the general case $\bm{X}^{l-1}\in\mathbb{R}^{n\times p}$. This leads to an overestimation of the sensitivity and to unnecessarily large noise addition. We directly analyze the change in $\bm{X}^{l-1}\in\mathbb{R}^{n\times p}$ and derive a tighter bound on the sensitivity $\Delta_{l}$. By using our proposed adjacency notion from Equation (12), we establish the following result (proof provided in Appendix A.1):
@@ 3.1.2 Improved sensitivity bound:
Let $\bm{A}^{\prime}$ be defined as in Equation (12), and consider the sensitivity $\Delta_{l}=\|\bm{A}^{\prime}\bm{X}^{l}-\bm{A}\bm{X}^{l}\|_{F}$. Then, $\Delta_{l}$ is upped-bounded as follows:
$\Delta_{l}\leq\max_{i}\|\bm{X}_{i,:}^{l}\|_{2}\triangleq\hat{\Delta}_{l}.$ (14)
This improved bound $\hat{\Delta}_{l}$ is generally tighter than the previous bound $\Delta_{l}^{\text{prior}}=\sqrt{p}\cdot\|\bm{X}_{l}\|_{\max}$, as it eliminates the dependence on $\sqrt{p}$ and directly relates to the maximum row norms of $\bm{X}^{l}$. By reducing the estimated sensitivity, we can add less noise while still satisfying Differential Privacy, thereby improving the utility of the algorithm.
3.1.3 Privacy proof:
To address these issues, we propose a result with a new proof of the Differential Privacy guarantees for our algorithm in Theorem 3.2, whose proof is in Appendix A.2.2. This proof clarifies the ambiguities or errors present in the previous works (see Appendix A.2), and ensures that our privacy guarantees are met.
@@ 3.1.3 Privacy proof:
Assuming $\delta\in(0,1)$ and $\epsilon>0$ such that $\delta\leq\exp{(-\frac{\epsilon}{4})}$, the Private Power Method with $\Delta_{l}=\max_{i}\|\bm{X}_{i:}^{l}\|_{2}$ satisfies ($\epsilon,\delta$)-Differential Privacy.
3.1.4 Improved convergence bound:
Theorem 3.3 (Bound for the noisy power method (NPM) (Balcan et al., 2016)).
Let reals $k\leq q\leq p$. Let $\bm{U}_{q}\in\mathbb{R}^{d\times q}$ be the top-$q$ eigenvectors of a positive semi-definite matrix $\bm{A}$ and let $\sigma_{1}\geq\cdots\geq\sigma_{d}\geq 0$ denote its eigenvalues and fix $\eta=O\left(\frac{\sigma_{q}}{\sigma_{k}}\cdot\min\left\{\frac{1}{\log\left(\frac{\sigma_{k}}{\sigma_{q}}\right)},\frac{1}{\log\left(\tau d\right)}\right\}\right)$. If at every iteration $l$ of the NPM $\bm{G}_{\ell}$ satisfies Assumption 2.3, then after
$L=\Theta\left(\frac{\sigma_{k}}{\sigma_{k}-\sigma_{q+1}}\log\left(\frac{\tau d}{\eta}\right)\right).$
$\|(\bm{I}-\bm{X}^{L}{\bm{X}^{L}}^{\top})\bm{U}_{k}\|_{2}\leq\eta.$ (15)
Theorem 3.4 (Improved PPM with runtime dependent bound).
Let $\bm{A}\in\mathbb{R}^{d\times d}$ be a symmetric data matrix. Fix target rank $k$, intermediate rank $q\geq k$ and iteration rank $p$ with $2q\leq p\leq d$. Suppose the number of iterations $L$ is set as $L=\Theta(\frac{\sigma_{k}}{\sigma_{k}-\sigma_{q+1}}\log(d))$. Let $\delta\in(0,1)$ and $\epsilon>0$ be privacy parameters such that $\delta\leq\exp{(-\frac{\epsilon}{4})}$
Then with approximate DP parameters ($\epsilon$,$\delta$) and assuming Assumption 2.3 is verified, Algorithm 1 with $\Delta_{l}=\max_{i}\|\bm{X}_{i:}^{l}\|_{2}$ gives us the following runtime-dependent convergence bound:
$\eta=O\left(\frac{\epsilon^{-1}\max_{i,l}\|\bm{X}_{i:}^{l}\|_{2}\sqrt{Ld\log(1/\delta)\log(L)}}{\sigma_{k}-\sigma_{q+1}}\right).$ (16)
Note: This bound is at least as tight as the one proposed in (Balcan et al., 2016), and can be in practice much tighter depending on $\bm{A}$. Our privacy proof (given in Theorem 3.2) does not restrict us to have $\epsilon\leq 1$. We provide the proof in Appendix A.2.2.
4 Analysis of Runtime-independent bounds
4.1 Proposed runtime-independent convergence bound
We presented in Theorem 3.4 a runtime-dependent convergence bound involving $\max_{i}\|\bm{X}_{i,:}\|_{2}$, which is only known during the execution of the algorithm. To provide a more practical analysis, we now derive a runtime-independent convergence bound in Theorem 4.1. This bound is a combination of two cases:
  • High Coherence Case: Applies to matrices with large coherence parameters, and the bound is independent of the matrix’s coherence.
  • Low Coherence Case: Applies to matrices with low $\mu_{1}$-coherence, with the bound depending explicitly on $\mu_{1}(\bm{A})$.
We demonstrate that our method achieves a convergence bound that is at least as tight as both of these bounds.
We note that for a matrix $\bm{A}$, the proposed bound is tighter than the one proposed in (Balcan et al., 2016) when $\min(1,\mu_{1}(\bm{A}))\leq\sqrt{p\cdot\log({d})}\cdot\mu_{0}(\bm{A})$. This is likely to happen in practice, as we highlight in Section 6.1.1.
@@ 4.1 Proposed runtime-independent convergence bound
Improved PPM with runtime independent bound
Let $\bm{A}\in\mathbb{R}^{d\times d}$ be a symmetric data matrix. Fix target rank $k$, intermediate rank $q\geq k$ and iteration rank $p$ with $2q\leq p\leq d$. Suppose the number of iterations $L$ is set as $L=\Theta(\frac{\sigma_{k}}{\sigma_{k}-\sigma_{q+1}}\log(d))$. Let $\delta\in(0,1)$ and $\epsilon>0$ be privacy parameters such that $\delta\leq\exp{(-\frac{\epsilon}{4})}$ Then with approximate DP parameters ($\epsilon$,$\delta$) and assuming Assumption 2.3 is verified, Algorithm 1 gives us the following runtime-independent convergence bound:
$\eta=O\left(\frac{\epsilon^{-1}\cdot\min(1,\mu_{1}(\bm{A}))\cdot\sqrt{Ld\log(1/\delta)\log(L)}}{\sigma_{k}-\sigma_{q+1}}\right).$ (17)
@@ 5 Decentralized version
Table 1: Overhead of using Secure Aggregation in our protocol
Client Server
Communication $O(log(s)+n^{2})$ $O(s\cdot log(s)+sn^{2})$
Computation $O(log^{2}(s)+n^{2}\cdot log(s))$ $O(s\cdot log^{2}(s)+sn^{2}\cdot log(s))$
@@
In this section, we consider a decentralized setting where the matrix $\bm{A}$ is distributed across multiple clients. Specifically, each client $i$ holds a private matrix $\bm{A}^{(i)}\in\mathbb{R}^{n\times n}$, such that the global matrix is the sum of these local matrices: $\bm{A}=\sum_{i=1}^{s}\bm{A}^{(i)}.$
The goal is for the clients to collaboratively compute an orthonormal basis for the range of $\bm{A}$, similar to the centralized randomized power method, but without revealing their individual private matrices $\bm{A}^{(i)}$ to the server or to other clients.
@@
The randomized power method involves linear operations, making it well-suited for parallelization and distributed computation. Balcan et al. (2016) proposed a private and federated power method using communication over public channels between clients and a server. However, this approach relies on Differential Privacy, since the data exchanged can be observed by everyone, which requires high levels of noise to ensure privacy.
@@
To enhance privacy while retaining the benefits of distributed computation, we use Secure Aggregation, a lightweight Multi-Party Computation protocol. Secure Aggregation allows clients to collaboratively compute sums without revealing individual data, enabling the use of distributed DP. This approach offers privacy guarantees similar to central DP and eliminates the need for a trusted curator. Distributed DP has been extensively studied in the literature (Chen et al., 2021; Kairouz et al., 2021; Goryczka et al., 2013; Wei et al., 2024; Ghazi et al., 2019).
@@
In distributed DP, each client adds carefully calibrated noise to their local contributions before participating in the Secure Aggregation protocol. The noise is designed such that the sum across clients has a variance comparable to that used in central DP, thereby achieving similar Differential Privacy guarantees without requiring a trusted aggregator, and allowing clients to keep the data locally.
@@
We operate under the “honest-but-curious“ threat model, where clients follow the protocol correctly but may attempt to learn additional information from received data. We also assume there are no corrupted or dropout users during the computation. For simplicity, this paper focuses on a basic implementation of distributed DP, neglecting the effects of data quantization or errors introduced by modular arithmetic. We refer to Kairouz et al. (2021) for an implementation of Distributed DP that takes these effects into account.
@@
Finally, it is possible to use a fully decentralized and differentially private protocol to perform averaging without relying on a central server for Secure Aggregation, as demonstrated by Sabater et al. (2022). Adopting such approaches can further enhance decentralization while maintaining similar utility and communication costs.

### v3

@@ 2.3 Existing convergence bounds
To our knowledge, the strongest convergence bound for the privacy-preserving randomized power method (Algorithm 1) uses the bound $\Delta_{l}\leq\Delta_{l}^{prior}$ and is given by Balcan et al. (2016) in their Corollary 3.1. The following theorems rely on the conditions in Assumption 2.1, which are enforced by judicious choice of the added noise.
@@
Let $\bm{A}\in\mathbb{R}^{n\times n}$ be a symmetric matrix. Fix a target rank $k$, an intermediate rank $q\geq k$, and an iteration rank $p$, with $k\leq q\leq p$. Let $\bm{U}_{q}\in\mathbb{R}^{n\times q}$ be the top-$q$ eigenvectors of $\bm{A}$ and let $\lambda_{1}\geq\cdots\geq\lambda_{n}\geq 0$ denote its eigenvalues. Let us fix $\eta=O\left(\frac{\lambda_{q}}{\lambda_{k}}\cdot\min\left\{\frac{1}{\log\left(\frac{\lambda_{k}}{\lambda_{q}}\right)},\frac{1}{\log\left(\tau n\right)}\right\}\right).$
@@
$\|\bm{G}_{\ell}\|_{2}$ $=O\left(\eta(\lambda_{k}-\lambda_{q+1})\right)\,,\quad\text{and}\quad\|\bm{U}_{q}^{\top}\bm{G}_{\ell}\|_{2}=O\left(\eta\left(\lambda_{k}-\lambda_{q+1}\right)\frac{\sqrt{p}-\sqrt{q-1}}{\tau\sqrt{n}}\right)\,.$ (4)
We now restate the Private Power Method major result in (Balcan et al., 2016).
Theorem 2.2 (Private Power Method (PPM), reduction to $s=1$ from the proof in Appendix C.1 from Balcan et al. (2016)).
Let $\bm{A}\in\mathbb{R}^{n\times n}$ be a symmetric data matrix. Fix target rank $k$, intermediate rank $q\geq k$, and iteration rank $p$, with $2q\leq p\leq n$. Suppose the number of iterations $L$ is set as $L=\Theta(\frac{\lambda_{k}}{\lambda_{k}-\lambda_{q+1}}\log(n))$. Let $\epsilon,\delta\in(0,1)$ be the differential privacy parameters. Let $\bm{U}_{q}\in\mathbb{R}^{n\times q}$ be the top-$q$ eigenvectors of $\bm{A}$ and let $\lambda_{1}\geq\cdots\geq\lambda_{n}\geq 0$ denote its eigenvalues. Then Algorithm 1 with $\Delta_{l}=\Delta_{l}^{prior}$ is ($\epsilon$, $\delta$)-DP and with probability at least 0.9
$\|(\bm{I}-\bm{X}^{L}({\bm{X}^{L}})^{\top})\bm{U}_{k}\|_{2}\leq\eta\quad\text{and}\quad\|(\bm{I}-\bm{X}^{L}({\bm{X}^{L}})^{\top})\bm{A}\|_{2}^{2}\leq\lambda_{k+1}^{2}+\eta^{2}\lambda_{k}^{2}$ (5)
[…]
3 Proposed Differentially Private Power Method Convergence Bounds
In this section, we present an enhanced Differentially Private Power Method. First, we introduce a generalized definition of adjacency that goes beyond single-entry changes in a PSD matrix to allow for more applications. Then, we modify the Private Power Method to calibrate DP noise using a new, strictly tighter sensitivity bound, eliminating the $\sqrt{p}$ (associated to target rank) factor in prior work. We finish by showing analytically that this refinement yields strictly sharper convergence guarantees.
Adjacency notion.
In prior works (Hardt and Price, 2014; Balcan et al., 2016), two datasets represented by PSD matrices $\bm{A}\in\mathbb{R}^{n\times n}$ and $\bm{A}^{\prime}\in\mathbb{R}^{n\times n}$ are considered adjacent (denoted $\bm{A}\sim\bm{A}^{\prime}$) if they differ by a change in a single element, with a Frobenius norm of at most 1:
@@ 3 Proposed Differentially Private Power Method Convergence Bounds
where $\bm{e}_{i},\bm{e}_{j}\in\mathbb{R}^{n}$ are canonical basis vectors, $c\leq 1$ represents the magnitude of the change, and $0\leq i,j<n$. This adjacency notion models a sensitive change as a modification to a single element in $\bm{A}$ to protect individual element-wise updates under Differential Privacy. However, since $\bm{A}$ is symmetric positive semi-definite and changes must preserve this property, this formulation restricts updates to be diagonal, hindering possible applications.
We propose a new, more general, notion of adjacency to allow for other types of updates:
@@
where $\bm{C}\in\mathbb{R}^{n\times n}$ is a symmetric matrix representing the update, subject to $\sqrt{\sum_{i=1}^{n}\|\bm{C}_{i,:}\|_{1}^{2}}\leq 1$.
The proposed adjacency notion is strictly more general than Equation (6). For example, setting $\bm{C}=c\cdot\bm{e}_{i}\bm{e}_{j}^{\top}$ recovers the original definition. $\bm{C}$ can have non-zero entries on the diagonal, anti-diagonal, or any symmetric pattern, allowing for a variety of updates maintaining symmetry.
In the context of recommender systems, where $\bm{A}=\bm{R}^{\top}\bm{R}$ represents the item-item similarity matrix and $\bm{R}\in\mathbb{R}^{m\times n}$ is the user-item interaction matrix, our proposed notion of adjacency enables element-wise modifications in $\bm{R}$ (i.e., protecting individual user-item interactions). Such changes in $\bm{R}$ propagate to multiple elements of $\bm{A}$, which could not be adequately accounted for under the previous adjacency definition (Equation (6)). By adopting our more general adjacency definition, we make our privacy guarantees applicable to a wider range of real-world scenarios.
Sensitivity bound.
The previous sensitivity bound $\Delta_{l}^{\text{prior}}$ (Hardt and Price, 2014; Balcan et al., 2016) for $\bm{A}\bm{X}^{l-1}$ defined in Equation (3) was estimated by extrapolating from the case where $\bm{X}^{l-1}\in\mathbb{R}^{n\times 1}$ to the general case $\bm{X}^{l-1}\in\mathbb{R}^{n\times p}$, leading to a dependence on $\sqrt{p}$. This leads to an overestimation of the sensitivity and to unnecessarily large noise addition. We directly analyze the change in $\bm{X}^{l-1}\in\mathbb{R}^{n\times p}$ and derive a strictly tighter bound on the sensitivity $\Delta_{l}$. By using our adjacency notion from Equation (7), we establish the following result (proof in Appendix A.1):
@@
Let $\bm{A}^{\prime}$ be defined as in Equation (7), and consider the sensitivity $\Delta_{l}=\sup_{\bm{A}\sim\bm{A}^{\prime}}\|\bm{A}^{\prime}\bm{X}^{l}-\bm{A}\bm{X}^{l}\|_{F}$. Then,
$\Delta_{l}\leq\max_{i}\|\bm{X}_{i:}^{l}\|_{2}\triangleq\hat{\Delta}_{l}.$ (8)
Privacy proof.
Our algorithm achieves differential privacy by iteratively adding calibrated noise at each round of the power method. The total privacy guarantee across iterations is then derived using results from adaptive composition of DP mechanisms, as initially proposed in (Bun and Steinke, 2016), allowing us to precisely quantify the cumulative privacy loss across multiple iterative steps. To clarify the ambiguities or errors present in the previous works (see Appendix A.2), and ensure that our privacy guarantees are met, we propose a result with a new proof of the Differential Privacy guarantees for our overall algorithm in Theorem 3.2, whose proof is in Appendix A.2.3.
@@
Let $\delta\in(0,1)$ and $\epsilon>0$ such that $\delta\leq\exp{(-\frac{\epsilon}{4})}$. Then, Algorithm 1 with $\Delta_{l}=\max_{i}\|\bm{X}_{i:}^{l}\|_{2}$ is ($\epsilon,\delta$)-Differentially Private.
Improved convergence bound.
Building on Equation (8), we present in Theorem 3.3 a strictly tighter convergence bound than the one proposed in (Balcan et al., 2016). Additionally, unlike past proofs (Hardt and Price, 2014; Balcan et al., 2016), our proposed privacy proof (given in Theorem 3.2) does not restrict $\epsilon\leq 1$. We provide the proof in Appendix A.3.
Theorem 3.3 (Improved PPM with Runtime-Dependent Bound).
Let $\bm{A}\in\mathbb{R}^{n\times n}$ be a symmetric data matrix. Fix target rank $k$, intermediate rank $q\geq k$ and iteration rank $p$ with $2q\leq p\leq n$. Suppose the number of iterations $L=\Theta(\frac{\lambda_{k}}{\lambda_{k}-\lambda_{q+1}}\log(n))$. Let $\delta\in(0,1)$ and $\epsilon>0$ be privacy parameters such that $\delta\leq\exp{(-\frac{\epsilon}{4})}$. Let $\bm{U}_{k}\in\mathbb{R}^{n\times k}$ be the top-$k$ eigenvectors of $\bm{A}$ and let $\lambda_{1}\geq\cdots\geq\lambda_{n}\geq 0$ denote its eigenvalues. Then Algorithm 1 is ($\epsilon$,$\delta$)-DP with $\Delta_{l}=\max_{i}\|\bm{X}_{i:}^{l}\|_{2}$ and with probability at least 0.9
$\|(\bm{I}-\bm{X}^{L}({\bm{X}^{L}})^{\top})\bm{U}_{k}\|_{2}\leq\eta\quad\text{and}\quad\|(\bm{I}-\bm{X}^{L}({\bm{X}^{L}})^{\top})\bm{A}\|_{2}^{2}\leq\lambda_{k+1}^{2}+\eta^{2}\lambda_{k}^{2}$ (9)
with $\text{with }\quad\eta=O\left(\frac{\epsilon^{-1}\max_{i,l}\|\bm{X}_{i:}^{l}\|_{2}\sqrt{Ln\log(1/\delta)\log(L)}}{\lambda_{k}-\lambda_{q+1}}\right).$ (10)
4 Proposed Runtime-Independent Convergence Bound
We presented in Theorem 3.3 a convergence bound involving $\max_{i}\|\bm{X}_{i:}\|_{2}$, which is only observable during the execution of the algorithm. To provide a more practical analysis, we now derive a runtime-independent convergence bound in Theorem 4.1 by careful bounding of $\max_{i}\|\bm{X}_{i:}\|_{2}$. We provide a proof in Appendix A.4. This bound makes it possible to have a tight analysis in two regimes:
  • Small $\mu_{0}$-coherence, small $p$: Previously proposed in Balcan et al. (2016), this bound is useful in a regime with small $\mu_{0}$ when computing few eigenvectors, with a dependence on $\sqrt{p\log({n})}\cdot\mu_{0}(\bm{A})$.
  • Larger $\mu_{0}$ or $p$: We propose a new bound tailored for the multi-dimensional power method, depending on $\mu_{1}(\bm{A})$, with a reduced dependence on the number of eigenvectors $p$.
We note that we are likely to be in the second regime in practice, as we highlight in Section 6.1.
@@ 4 Proposed Runtime-Independent Convergence Bound
Improved PPM with Runtime-Independent Bound
Let $\bm{A}\in\mathbb{R}^{n\times n}$ be a symmetric data matrix. Fix target rank $k$, intermediate rank $q\geq k$ and iteration rank $p$ with $2q\leq p\leq n$. Suppose the number of iterations $L$ is set as $L=\Theta(\frac{\lambda_{k}}{\lambda_{k}-\lambda_{q+1}}\log(n))$. Let $\bm{U}_{q}\in\mathbb{R}^{n\times q}$ be the top-$q$ eigenvectors of $\bm{A}$ and let $\lambda_{1}\geq\cdots\geq\lambda_{n}\geq 0$ denote its eigenvalues. Let $\delta\in(0,1)$ and $\epsilon>0$ be privacy parameters such that $\delta\leq\exp{(-\frac{\epsilon}{4})}$ Then Algorithm 1 is ($\epsilon$,$\delta$)-DP and we have with probability at least 0.9
$\|(\bm{I}-\bm{X}^{L}({\bm{X}^{L}})^{\top})\bm{U}_{k}\|_{2}\leq\eta\quad\text{and}\quad\|(\bm{I}-\bm{X}^{L}({\bm{X}^{L}})^{\top})\bm{A}\|_{2}^{2}\leq\lambda_{k+1}^{2}+\eta^{2}\lambda_{k}^{2}$ (11)
with $\text{with }\quad\eta=O\left(\frac{\epsilon^{-1}\cdot\min(\mu_{0}(\bm{A})\sqrt{p\cdot\log({n})},\mu_{1}(\bm{A}))\cdot\sqrt{Ln\log(1/\delta)\log(L)}}{\lambda_{k}-\lambda_{q+1}}\right).$ (12)
@@ 5 Decentralized version
Table 1: Per-iteration overhead: Secure Aggregation vs. Federated Learning.
Secure Aggregation Federated Learning
Client Server Client Server
Communication $O\bigl{(}\log(s)+np\bigr{)}$ $O\bigl{(}s\,[\log(s)+np]\bigr{)}$ $O(np)$ $O\bigl{(}s\,np\bigr{)}$
Computation $O\bigl{(}\log^{2}(s)+np\log(s)\bigr{)}$ $O\bigl{(}s\,[\log^{2}(s)+np\log(s)]\bigr{)}$ $O(np)$ $O\bigl{(}s\,np\bigr{)}$
@@
In this section, we consider a decentralized setting where the matrix $\bm{A}$ is distributed across multiple clients. Specifically, each client $i$ holds a private matrix $\bm{A}^{(i)}\in\mathbb{R}^{n\times n}$, such that the global matrix is the sum of these local matrices: $\bm{A}=\sum_{i=1}^{s}\bm{A}^{(i)}$. The goal is for the clients to collaboratively compute an orthonormal basis for the range of $\bm{A}$, similar to the centralized randomized power method, but without revealing their individual private matrices $\bm{A}^{(i)}$ to the server or to other clients.
@@
The randomized power method involves linear operations, making it well-suited for parallelization and distributed computation. Balcan et al. (2016); Guo et al. (2024) proposed private and federated power methods using communication over public channels between clients and a server. However, these approaches rely on local Differential Privacy, since the data exchanged can be observed by everyone, which requires high levels of noise to ensure privacy.
@@
To enhance privacy while retaining the benefits of distributed computation, we use Secure Aggregation, a lightweight Multi-Party Computation protocol. Secure Aggregation allows clients to collaboratively compute sums without revealing individual data, enabling the use of distributed DP. This approach offers privacy guarantees similar to central DP and eliminates the need for a trusted curator. Distributed DP has been extensively studied in the literature (Goryczka et al., 2013; Ghazi et al., 2019; Kairouz et al., 2021; Chen et al., 2021; Wei et al., 2024).
@@
In distributed DP, each client adds carefully calibrated noise to their local contributions before participating in the Secure Aggregation protocol. The noise is designed such that the sum across clients has a variance comparable to that used in central DP, thereby achieving similar Differential Privacy guarantees without requiring a trusted aggregator, and allowing clients to keep data locally.
@@
We operate under the honest-but-curious threat model, where clients follow the protocol correctly but may attempt to learn information from received data. We also assume that there are no corrupted or dropout users during the computation. For simplicity, this papers neglects the effects of data quantization or errors introduced by modular arithmetic in distributed DP and refer to Kairouz et al. (2021) for an implementation taking this into account.
@@
It is also possible to use a fully decentralized DP protocol to perform secure averaging without relying on a central server, as demonstrated by Sabater et al. (2022). Adopting such approaches can further enhance decentralization while maintaining similar utility and communication costs for the method.